OpenStudy (anonymous):

Find the value of 2 times the square root of 3, all over the sqaure root of 10 in simplest form.

OpenStudy (anonymous):

\[\frac{ 2\sqrt{3} }{ \sqrt{10} }\] is this the question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Rationalize the denominator, by multiplying the numerator and denominator by the radical in the denominator: \[\frac{ 2\sqrt{3} *\sqrt{10}}{ \sqrt{10}*\sqrt{10} }\]

OpenStudy (anonymous):

This simplifies to: \[\frac{ 2\sqrt{30} }{ \sqrt{100} }\]

OpenStudy (anonymous):

What is the square root of 100?

OpenStudy (anonymous):

50? Im not quite sure...

OpenStudy (anonymous):

What number can you square to give you 100? or what number can u multiply by itself to give you 100

OpenStudy (anonymous):

10

OpenStudy (anonymous):

good so now we have:\[\frac{ 2\sqrt{30} }{ 10 }\] Can you simplify this and tell me what you got?

OpenStudy (anonymous):

2 sqrt 3 over 5?

OpenStudy (anonymous):

Nope you can only divide a radical by a radical. Therefore it is incorrect to divide 30 by 10. But you can simplify the numbers that are not radicals, by dividing 2 by 2 in the numerator and 10 by 2 in the denominator: You should get as your answer - \[\frac{ \sqrt{30} }{ 5 }\]

OpenStudy (anonymous):

Do you understand how I got this?

OpenStudy (anonymous):

Oh Okay... yeah i got it now, thanks!